Re: Shift report 20000809 00:00-08:00

From: Konstantin Olchanski (olchansk@ux1.phy.bnl.gov)
Date: Wed Aug 09 2000 - 15:50:56 EDT

  • Next message: http account: "Shift report 20000809 08:00-16:00"

    On Wed, Aug 09, 2000 at 08:11:14AM -0400, http account wrote:
    > ------------------------------------------------------------
    > 3. Around 1:50, noticed event rate = 0 in DAQ. Stopped the
    >    run. Got a lot of EventBuilder error. MC called to ask 
    >    why scalers were stuck. Reset the scalers in LumiScaler.
    >    Individual ZDC and BB were counting, the others remained
    >    at 0. Visual scalers in FEH were stuck too. Reset them 
    >    but remained at 0. Called Konstantin and Dana. 
    >    Dana came around 3:00. Cycled power of some crates NIM 
    >    bins and got scalers back. See logbook for details. 
    >    Got the DAQ work when only camac and fb4 partitions were
    >    selected => TPC and DC partitions screw the DAQ. 
    >    Sent an email to Bob Scheetz to check TPC readout.
    > ------------------------------------------------------------
    
    
    The problem was caused by the TPC trigger board lockup. It was hard
    to find but easy to fix- by cycling power on the Slow Controls VME crate.
    
    
    -- 
    Konstantin Olchanski
    Physics Department, Brookhaven National Laboratory, Long Island, New York
    olchansk@bnl.gov
    



    This archive was generated by hypermail 2b29 : Wed Aug 09 2000 - 15:54:32 EDT